OpenStudy (anonymous):

Jules has 265$ at the begging of the month. What amount ,y , does he have left after x days if he spends 8.50 a day ? a. Write an equation that expresses the relationship between x and y. B. find two ordered pairs in the solution set of the equation that you wrote in part a.

OpenStudy (anonymous):

A) y=265-8.50x where x=1, 2, 3,4, ...

OpenStudy (anonymous):

so what is B?

OpenStudy (anonymous):

B) there are as many ordered pairs as there are days in the given month, for example after 6 days J has y=265-8.50(6)=214 dollars; as an ordered pair (6, 214) after 8 days J has y=265-8.50(8)=197 dollars; as an ordered pair (8, 197) etc.

OpenStudy (anonymous):

thanks :D

OpenStudy (anonymous):

np